OpenStudy (anonymous):

A boat's crew rowed 13.5 miles downstream, with the current, in 1.5 hours. The return trip upstream, against the current, covered the same distance, but took 4.5 hours. Find the crew's rowing rate in still water and the rate of the current.

OpenStudy (anonymous):

The crew's rowing rate in still water is ? miles per hour? The rate of the current is ? miles per hour?

OpenStudy (anonymous):

Crew rowing rate be x current rate be y x+y=13.5/1.5=9 x-y=13.5/4.5=3 adding the two equations we get 2x=12 x=6 miles per hour(Crew Rowing rate) 6+y=9 y=3(rate of current)

OpenStudy (anonymous):

thank you